shototodorokii:

Mary, Peter, and Lucy were picking chestnuts. Mary picked twice as many chestnuts as Peter. Lucy picked 2 kg more than Peter. Together the three of them picked 26 kg of chestnuts. How many kilograms did each of them pick?

babygirlimarie:

Ok do you want the answer or just a way to figure it out ?

shototodorokii:

Just the answer

babygirlimarie:

OK

shototodorokii:

i already solved it

Zaiaz:

so you want to make sure that ur answer is correct?

shototodorokii:

no, i got the answer i just want to see if you guys get it correct

Zaiaz:

okay

shototodorokii:

I'm closing this question in 10 minutes

shototodorokii:

six minutes left

shototodorokii:

4 minutes left

shototodorokii:

if you still haven't answer when the time goes out, you could message me.

shototodorokii:

3 minutes left

shototodorokii:

2 minutes

shototodorokii:

1 minute

shototodorokii:

Time is up

shototodorokii:

turn in answers now

shototodorokii:

Let xx be the amount Peter picked. Then Mary and Lucy picked 2x2x and x+2x+2, respectively. So x+2x+x+2=26x+2x+x+2=26 4x=244x=24 x=6x=6 Therefore, Peter, Mary, and Lucy picked 6, 12, and 8 kg, respectively.

shototodorokii:

That is the answer

shototodorokii:

If you got it correct then good for you

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @shototodorokii Let xx be the amount Peter picked. Then Mary and Lucy picked 2x2x and x+2x+2, respectively. So x+2x+x+2=26x+2x+x+2=26 4x=244x=24 x=6x=6 Therefore, Peter, Mary, and Lucy picked 6, 12, and 8 kg, respectively. \(\color{#0cbb34}{\text{End of Quote}}\) Not tryna be mean buttt